OpenStudy (rayman5766):

Multiply. Express your answer in lowest terms. 1 2 - . - 3 9

OpenStudy (rayman5766):

@iGreen

OpenStudy (igreen):

Multiply the tops and the bottoms. \(\dfrac{1}{3} \cdot \dfrac{2}{9} = \dfrac{1 \cdot 2}{3 \cdot 9}\) What's 1 * 2 and what's 3 * 9? @rayman5766

OpenStudy (rayman5766):

2 and 27 I know I just was no sheer if it is a trick question

OpenStudy (igreen):

Yep, you got it. So we have \(\dfrac{2}{27}\), that's your answer..because you can't simplify it any more.

OpenStudy (rayman5766):

ok can u help me with other

OpenStudy (igreen):

Sure. Just close this one and open another.